if the perimeter is 184 and the length is 3 times as long as the width, what is the length and the width
Answer by ptaylor(2198)

Perimeter(P) =2L+2W
Length=3W
So:
184=2*3W+2W
184=8W
W=23 Units
L=3W=3*23=69 Units
CK
184=2*69+2*23
184=138+46
184=184
Hope this helps---ptaylor